OpenStudy (anonymous):

Which fraction is equivalent to 68%? A. 17/50 B. 17/25 C.34/25 D. 680/100

OpenStudy (anonymous):

SHOW WORK = MEDAL

OpenStudy (henrietepurina):

68% = 68/100 =34/50 =17/25 so B

OpenStudy (henrietepurina):

TaDa!